# Recorded Versus Synthetic Ground Motions: A Comparative Analysis of Structural Seismic Responses

### v1

@@ Abstract
This paper presents a comparative analysis of structural seismic responses under two types of ground motion inputs: (i) synthetic motions generated by stochastic ground motion models and (ii) recorded motions from an earthquake database. Five key seismic response metrics—probability distributions, statistical moments, correlations, tail indices, and variance-based global sensitivity indices—are systematically evaluated for two archetypal structures: a 12-story medium-period building and a high-rise long-period tower. Both ground motion datasets are calibrated to a shared response spectrum, ensuring consistency in spectral characteristics, including spectral median, variance, and correlation structure. The analysis incorporates both aleatory uncertainties from ground motion variability and epistemic uncertainties associated with structural parameters, providing a comprehensive comparison of seismic responses. The results demonstrate close agreement in global response characteristics, including distributions, correlations, and sensitivity indices, between synthetic and recorded motions, with differences typically within 15%. However, significant discrepancies are observed under extreme conditions, particularly in tail behavior, higher-order moments, and drift responses of long-period structures, with differences exceeding 50%. These discrepancies are attributed to the non-Gaussian features and complex characteristics inherent in recorded motions, which are less pronounced in synthetic datasets. The findings support the use of synthetic ground motions for evaluating global seismic response characteristics, while highlighting their limitations in capturing rare-event behavior and long-period structural dynamics.
@@ 1 Introduction
Quantifying the variability in seismic responses of structures subjected to ground motion inputs is a fundamental challenge in performance-based earthquake engineering (PBEE) [1, 2, 3, 4, 5] and seismic risk assessment [6, 7, 8, 9]. Structural responses to seismic excitations are primarily governed by two sources of input uncertainty: variability in ground motion characteristics (aleatory uncertainty) and uncertainties in structural parameters (epistemic uncertainty). A comprehensive understanding of stochastic structural responses is crucial for enhancing the reliability of structural designs and accurately assessing seismic risks, particularly under extreme scenarios involving limit state exceedance. Evaluating different seismic response metrics enables systematic comparison of uncertainties across diverse scenarios, supporting risk-informed structural design.
The current practice for defining ground motion inputs typically involves generating a target response spectrum to characterize the seismic hazard at a site. The uniform hazard spectrum (UHS), derived from probabilistic seismic hazard analysis, provides spectral accelerations at various periods corresponding to a constant exceedance probability (e.g., 2% probability of exceedance in 50 years). While UHS remains widely used, it tends to be overly conservative because it does not capture the correlations between spectral accelerations at different periods. To address this, the conditional mean spectrum (CMS) has been developed as a more realistic alternative by conditioning the spectral accelerations at multiple periods on a target period of interest, thereby capturing these correlations and avoiding unnecessary conservatism [10]. In addition, ground motion models (GMMs) offer a site-specific approach, generating or selecting spectra based on parameters such as magnitude, distance, and local soil conditions [11, 12]. These methodologies, encompassing both recorded and synthesized signals generated from physics-based [13, 12] or statistics-based models [14, 15], offer varying degrees of fidelity and computational efficiency, thereby supporting refined ground motion selection or generation tailored to PBEE and risk assessment applications.
Provided with a target response spectrum, PBEE often requires selecting or generating spectrum-consistent ground motions. Ground motions are typically categorized into two types: recorded motions obtained from past earthquake events and synthetic motions generated using stochastic ground motion models (SGMMs). Recorded motions are widely regarded as realistic representations of earthquake phenomena, making them a preferred choice in design contexts due to their direct connection to observed seismic events [16, 17, 18]. However, their availability is often constrained, especially for rare events or specific seismic scenarios, and they may not encompass the full spectrum of potential earthquake characteristics. In contrast, synthetic motions offer high flexibility, allowing the simulation of a broad array of seismic scenarios with controlled spectral characteristics [14, 19, 20, 21]. This flexibility makes synthetic motions invaluable for seismic risk assessment. Despite their utility, synthetic motions are often criticized for their inability to replicate the complex, non-Gaussian features inherent in real ground motions, such as near-fault pulse-like effects and extreme tail behaviors [16, 22].
@@
A comparison between recorded and synthetic ground motions is essential to evaluate their respective strengths and limitations, particularly in the context of various structural response metrics. While previous studies [23, 24, 25, 26, 27, 28] have explored specific metrics, such as peak response distributions or spectral accelerations, comprehensive evaluations spanning a broader range of stochastic response metrics remain limited. Moreover, as seismic design increasingly integrates uncertainty quantification (UQ), variance-based sensitivity analysis has emerged as a robust tool for identifying key contributors to response variability. This enables engineers to prioritize critical uncertainties while simplifying those with lesser impacts on seismic performance. However, comparative analyses of seismic responses from the perspective of UQ [29, 30, 31] and variance-based sensitivity analysis [32, 33, 34] are underexplored. Addressing this gap, the present study systematically compares stochastic seismic responses for synthetic and recorded ground motions derived from a shared target response spectrum, ensuring consistency in spectral characteristics.
@@
This study evaluates five key stochastic seismic response metrics: probability distributions, statistical moments and normalized moments (mean, variance, skewness, kurtosis), correlations, tail indices, and variance-based sensitivity indices. Both synthetic and recorded ground motion datasets are calibrated to an equivalent target spectrum, maintaining alignment in spectral median, variance, and correlation characteristics. The sensitivity analysis incorporates both ground motion variability and structural parameter variability, complemented by a parametric study exploring the effects of input parameter distributions on sensitivity patterns. These metrics collectively capture various aspects of seismic response variability, ranging from global characteristics (e.g., distributions and correlations) to extreme behaviors (e.g., tail indices and higher-order moments) and sensitivity patterns (e.g., the relative influence of aleatory and epistemic uncertainties). Two archetypal structures are analyzed: a 12-story medium-period building and a 110-story long-period tower. These structures represent distinct dynamic regimes, facilitating an evaluation of how ground motion characteristics influence structural responses across varying periods. By systematically comparing these metrics, the study provides critical insights into the strengths and limitations of synthetic and recorded ground motions in the context of seismic design and risk assessment.
The findings demonstrate consistent agreement in global response characteristics, such as distributions, statistical moments, and correlations, between synthetic and recorded ground motions, with differences typically within 15%. However, significant discrepancies emerge under extreme conditions, exceeding 50%, particularly for higher-order moments, tail indices, and long-period structural responses. These differences are attributed to the non-Gaussian features of recorded motions, which are less pronounced in synthetic motions constrained by Gaussian assumptions. Additionally, the differences in seismic performance can be linked to limitations in ground motion selection and generation methods, including insufficient spectral matching and inherent variability in mapping the spectral acceleration to the ground motion time histories. The comparison confirms the effectiveness of synthetic motions calibrated to target spectra, particularly for global response characteristics, while emphasizing the need for caution in scenarios involving rare events and long-period structures.
@@
This paper is organized as follows: Section 2 introduces the five stochastic seismic response metrics, the ground motion datasets, and the two structural archetypes. Section 3 presents a comparative evaluation of seismic responses for both ground motion types across the two structures, highlighting key trends and insights. Section 4 synthesizes the results and summarizes the main findings. Finally, Section 5 discusses the implications of the findings for seismic design and risk assessment.
@@ 2.2 Metrics for stochastic seismic responses
    where $M_{k}$ ($k=1,...,4$) represents the mean, variance, skewness, and kurtosis, respectively.
@@
    Tail indices: Tail indices characterize the extremity of response distributions, which is critical for evaluating rare-event probabilities. The Hill’s estimator [35] is adopted to quantify tail behavior:
@@
    Variance-based sensitivity indices: Sobol’ sensitivity indices [32, 36] are adopted to quantify the contributions of ground motion variability and structural parameter uncertainty to response variability. The group-wise Sobol’ index measures the proportion of output variability attributable to a group of input variables $\boldsymbol{X}_{\mathbf{u}}$:
@@ 2.3 Synthetic and recorded ground motion datasets
Two ground motion datasets—synthetic and recorded—are developed to facilitate a comparative analysis of seismic responses. The GMM-based unconditional target spectra is adopted in this study. Both datasets are calibrated to a common target spectrum derived from the Ground Motion Prediction Equation (GMPE) by [11] and the spectral correlation model by [37]. This target spectrum represents a moderate seismic hazard scenario, serving as a benchmark for ensuring consistency between the datasets. Table 1 summarizes the seismic hazard parameters defining the target spectrum.
The recorded dataset consists of 2,000 ground motions selected from the PEER Next Generation Attenuation (NGA)-West 2 database [38]. The synthetic dataset comprises 2,000 ground motions generated using a spectral-compatibility method [19], calibrated to match the statistical characteristics of the target spectrum. Ground motion selection and generation follow standardized procedures [17, 19], as detailed in A. Each ground motion is treated as a discretized realization of the ground motion stochastic process, which is used in the comparative analysis.
@@
Figure 3: Spectral correlations of (a) the target spectrum, (b) recorded motions, and (c) synthetic motions. Contours represent empirical correlation coefficients versus $T_{1}$ and $T_{2}$, computed using 2,000 ground motions in each dataset.
2.4 Structural archetypes
Two structural archetypes, representing medium-period and long-period structures, are employed to investigate the influence of ground motion characteristics on structural responses across distinct dynamic regimes:
@@ 2.4 Structural archetypes
    Medium-period building: This archetype is a 12-story structure with a fundamental period of 1.09 seconds. This structure represents a typical mid-rise multi-degree-of-freedom (MDOF) building, whose seismic responses are dominated by lower modes.
@@
    Long-period tower: A 110-story structure with a fundamental period of 9.69 seconds, representing a high-rise MDOF tower. Its seismic responses are significantly influenced by higher modes and low-frequency ground motion components.
Both archetypes are modeled using nonlinear finite element analysis in OpenSees [39]. Nonlinear lateral force–displacement behavior is represented using bilinear springs, with a uniform story height of 3.65 meters. The seismic responses of interest include peak peak interstory drift ratios (IDRs) and peak floor accelerations (PFAs), which are widely used in seismic performance evaluation.
Uncertainties in structural parameters $\boldsymbol{X}_{\mathbf{S}}$, including damping ratio, floor weight, yield force, story stiffness, and strength hardening ratio, are modeled using standard distribution models. Table 2 summarizes these distributions across four cases: Case 1 assumes uniform distributions, while Cases 2–4 employ lognormal and truncated Gaussian distributions with decreasing coefficients of variation (c.o.v). These uncertainties are consistently applied to both structural archetypes in the comparative analysis. Figure 4 illustrates the structural models and the probability density functions (PDFs) of the parameters. Note that four test cases with different distribution models are considered to ensure the robustness of the comparative study conclusions.
Figure 4: Structural archetypes: (a) medium-period building, (b) long-period tower, and (c) PDFs of uncertain structural parameters across four cases.
Table 2: Probabilistic distributions of uncertain structural parameters for the structural archetypes.
@@ 3 Comparative analysis of seismic responses: synthetic versus recorded motions
This section presents a detailed comparative evaluation of seismic response metrics for synthetic and recorded ground motions. Five key metrics are analyzed: distributions, statistical moments, correlations, tail indices, and sensitivity indices. The first four metrics focus exclusively on aleatory uncertainties by considering ground motion variability, while structural parameters are fixed at their mean values, as specified in Case 1 of Table 2, ensuring that the analysis isolates the effects of ground motion variability on seismic responses. The fifth metric incorporates both ground motion variability and structural parameter uncertainty to evaluate their combined effect and relative contributions to seismic response variability. The analysis is conducted for both structural archetypes, examining trends in global and extreme behaviors.

### v2

# Recorded Versus Synthetic Spectral-compatible Ground Motions: A Comparative Analysis of Structural Seismic Responses

@@ Abstract
This paper presents a comparative analysis of structural seismic responses under two types of ground motion inputs: (i) synthetic motions generated by stochastic spectral-compatible ground motion models and (ii) recorded motions from an earthquake database. Both ground motion datasets are calibrated to a shared target response spectrum to ensure consistent spectral median, variance, and correlation structure. Five key stochastic response metrics—probability distributions, statistical moments, correlations, tail indices, and variance-based global sensitivity indices—are systematically evaluated for two representative structures: a medium-period building and a limiting case of a long-period tower. The comparison accounts for uncertainties both from ground motion and structural parameters. The results reveal that synthetic motions closely replicate recorded motions in terms of global response behavior—including distributions, mean and variance, correlation structure, and dominant uncertainty sources—indicating their suitability for routine seismic design and parametric studies. However, substantial differences emerge in response extremes for long-period structures, particularly in metrics governed by rare events, such as higher-order moments and tail behavior. These differences, which often exceed 50%, can be attributed to the non-Gaussian features and complex characteristics inherent in recorded motions, which are less pronounced in synthetic datasets. The findings support the use of synthetic ground motions for evaluating global seismic response characteristics, while highlighting their limitations in capturing rare-event behavior and long-period structural dynamics.
@@ 1 Introduction
Quantifying the variability in seismic responses of structures subjected to ground motion inputs is a fundamental challenge in performance-based earthquake engineering (PBEE) moehle2004framework , broccardo2016application , iervolino2017assessing , o2019conceptual , zhong2023surrogate and seismic risk assessment ellingwood2009quantifying , jalayer2010structural , taflanidis2011simulation , kim2025efficient . Structural responses to seismic excitations are primarily governed by two sources of input uncertainty: variability in ground motion characteristics and uncertainties in structural parameters. A comprehensive understanding of stochastic structural responses is crucial for enhancing the reliability of structural designs and accurately assessing seismic risks, particularly under extreme scenarios involving limit state exceedance. Evaluating different seismic response metrics enables systematic comparison of uncertainties across diverse scenarios, supporting risk-informed structural design.
The current practice for defining ground motion inputs typically involves generating a target response spectrum to characterize the seismic hazard at a site. The uniform hazard spectrum (UHS), derived from probabilistic seismic hazard analysis, provides spectral accelerations at various periods corresponding to a constant exceedance probability (e.g., 2% probability of exceedance in 50 years). However, UHS does not represent the response spectrum of any single ground motion and is generally conservative. This limitation has led to the development of advanced spectra that incorporate additional probabilistic information, such as spectral acceleration variability wang_ground_2011 and spectral correlation across distinct periods baker2008correlation . An advanced spectrum containing a complete probability distribution is termed an “unconditional spectrum” baker2018improved . When response spectra follow a joint log-normal distribution, the unconditional spectrum is fully defined by the mean, variance, and correlation of spectral accelerations across a specific range of periods. Using the properties of the joint lognormal distribution, Baker baker2011conditional developed the conditional mean spectrum (CMS), which refines spectral shapes by conditioning spectral accelerations at multiple periods on a target period of interest baker2011conditional .
Provided with a target response spectrum, PBEE often requires selecting or generating spectrum-consistent ground motions. Ground motions are typically categorized into two types: recorded motions obtained from past earthquake events and synthetically generated ground motions. Recorded motions are widely regarded as realistic representations of earthquake phenomena, making them a preferred choice in design contexts due to their direct connection to observed seismic events katsanos2010selection , baker2018improved , jalayer2020intensity . However, their availability is often constrained, especially for rare events or specific seismic scenarios, and they may not encompass the full spectrum of potential earthquake characteristics. In contrast, synthesized ground motions can be generated from physics-based approaches (e.g., taborda2015physics , boore2015revisions ) or classical statistics-based approaches (e.g., modulated filtered white noise models rezaeian2010simulation , su2024review , spectral-compatible methods gupta_synthesizing_1993 , cacciola_generation_2012 , zentner_procedure_2014 , yanni2024probabilistic and many others mobarakeh_simulation_2002 , yamamoto_stochastic_2013 ). More recently, deep generative models and physics-informed operator learning frameworks aquib2024broadband , esfahani2023tfcgan , lehmann2025multiple have received wide attentions, which complement traditional methods by enabling more flexible and data-driven modeling of complex seismic phenomena. Synthetic motions offer high flexibility, allowing the simulation of a broad array of seismic scenarios with controlled spectral characteristics kim2024adaptive , su2024importance , mobarakeh_simulation_2002 . Among these synthetic approaches, spectral-compatible stochastic ground motion models (SGMMs) gupta_synthesizing_1993 , cacciola_generation_2012 , zentner_procedure_2014 , yanni2024probabilistic are commonly used when recorded data are insufficient, yet their reliability in predicting structural response remains a key concern katsanos2010selection , kwong2015selection . This study evaluates spectral-compatible approaches by comparing recorded and synthetic motions with similar spectral characteristics, including spectral median, variance, and correlation structure.
A comparison between spectral-compatible recorded and synthetic ground motions is essential to evaluate their respective strengths and limitations, particularly in the context of various structural response metrics. A recent study rezaeian2024findings offers a comprehensive review of ground motion model validation, classifying validation metrics into two broad categories: ground motion characteristics and structural responses. Ground motion characteristics encompass waveform properties, response spectral intensity measures, peak parameters, and other intensity measures. In contrast, structural response metrics focus on the behavior of both idealized burks2014validation , galasso2020validation and realistic structural models fayaz2021methodology , munjy2022validation , considering quantities such as maximum inter-story drift ratios, peak floor accelerations, and energy-based measures chandramohan2016quantifying , tsioulou2019validation , karimzadeh2019use . Quantitative statistics—including median, dispersion, correlation, and kurtosis—can be evaluated for any scalar validation metric using suites of historical earthquake records alongside comparable synthetic simulations.
@@
While previous studies galasso2013validation , pei2014variability , karimzadeh2017assessment , li2021evaluation have explored specific metrics, such as peak response distributions or spectral accelerations, comprehensive evaluations spanning a broader range of stochastic response metrics remain limited. Moreover, as seismic design increasingly integrates uncertainty quantification (UQ), variance-based sensitivity analysis has emerged as a robust tool for identifying key contributors to response variability sobol2001global , padgett2007sensitivity , lamprou2013life . This enables engineers to prioritize critical uncertainties while simplifying those with lesser impacts on seismic performance. Despite its potential, limited research has applied variance-based sensitivity analysis to compare the influence of different uncertainty sources on structural responses under synthetic versus recorded excitations erazo2016uncertainty , kim2025uncertainty , smith2024uncertainty .
In this context, several prior studies have considered the effects of matched response spectra on structural response but often with limited scope. For example, Bijelic et al. bijelic2014seismic evaluated the seismic response of a tall building under recorded and simulated motions, while Iervolino et al. iervolino2010spectral assessed the role of spectral shape using artificial and real ground motions for single-degree-of-freedom systems. More recently, Bijelic et al. bijelic2018validation examined fragility-based collapse performance for various structures using physics-based synthetic motions, but with a limited number of records and without explicit analysis of spectral correlation. Notably, these studies did not match full spectral correlation matrices across periods. To our knowledge, no prior study has simultaneously matched and evaluated spectral median, variance, and correlation in synthetic and recorded datasets, nor assessed their impact across a comprehensive range of stochastic structural response metrics.
@@
This study aims to fill this gap by systematically comparing stochastic seismic response metrics between synthetic and recorded ground motions that are calibrated to the same unconditional target spectrum. Specifically, five response metrics are investigated: probability distributions, statistical moments and normalized moments (mean, variance, skewness, kurtosis), correlations, tail indices, and variance-based sensitivity indices. Both synthetic and recorded ground motion datasets are calibrated to an equivalent target spectrum, maintaining alignment in spectral median, variance, and correlation characteristics. The sensitivity analysis incorporates both ground motion variability and structural parameter variability, complemented by a parametric study exploring the effects of input parameter distributions on sensitivity patterns. These metrics collectively capture various aspects of seismic response variability, ranging from global characteristics (e.g., distributions and correlations) to extreme behaviors (e.g., tail indices and higher-order moments) and sensitivity patterns (e.g., the relative influence of uncertainties between ground motions and structures). Two archetypal structures are analyzed: a 12-story medium-period building and a 110-story long-period tower. These structures represent distinct dynamic regimes, facilitating an evaluation of how ground motion characteristics influence structural responses across varying periods. By systematically comparing these metrics, the study provides critical insights into the strengths and limitations of synthetic and recorded ground motions in the context of seismic design and risk assessment.
@@
This paper is organized as follows: Section 2 introduces the five stochastic seismic response metrics, the ground motion datasets, and the two structural models. Section 3 presents a comparative evaluation of seismic responses for both ground motion types across the two structures, highlighting key trends and insights. Section 4 synthesizes the results and summarizes the main findings. Finally, Section 5 discusses the implications of the findings for seismic design and risk assessment.
@@ 2.2 Metrics for stochastic seismic responses
    where $M_{k}$ (k=1,,4k=1,...,4) represents the mean, variance, skewness, and kurtosis, respectively.
@@
    Tail indices: Tail indices characterize the extremity of response distributions, which is critical for evaluating rare-event probabilities. The Hill’s estimator huisman2001tail is adopted to quantify tail behavior:
@@
    Variance-based sensitivity indices: Sobol’ sensitivity indices sobol2001global , saltelli2008global are adopted to quantify the contributions of ground motion variability and structural parameter uncertainty to response variability. The group-wise Sobol’ index measures the proportion of output variability attributable to a group of input variables $\boldsymbol{X}_{\mathbf{u}}$:
@@ 2.3 Synthetic and recorded ground motion datasets
Two ground motion datasets—synthetic and recorded—are developed to facilitate a comparative analysis of seismic responses. The GMM-based unconditional target spectra is adopted in this study. Both datasets are calibrated to a common target spectrum derived from the Ground Motion Prediction Equation (GMPE) by boore2014nga and the spectral correlation model by baker2008correlation . This target spectrum represents a moderate seismic hazard scenario, serving as a benchmark for ensuring consistency between the datasets. Table 1 summarizes the seismic hazard parameters defining the target spectrum.
The recorded dataset consists of 2,000 ground motions selected from the PEER Next Generation Attenuation (NGA)-West 2 database ancheta2014nga . The synthetic dataset comprises 2,000 ground motions generated using a spectral-compatibility method yanni2024probabilistic , calibrated to match the statistical characteristics of the target spectrum. Ground motion selection and generation follow standardized procedures baker2018improved , yanni2024probabilistic , as detailed in A. Each ground motion is treated as a discretized realization of the ground motion stochastic process, which is used in the comparative analysis.
It is important to note that the synthetic ground motions are generated using the evolutionary spectrum-based formulation of Yanni et al. (2024) yanni2024probabilistic , which employs a probabilistic time–frequency modulating envelope and an iterative spectral correction scheme. This variant does not rely on seed recorded accelerograms and is independent of the availability of historical earthquake records. Instead, it ensures consistent enforcement of the target spectral median, variance, and correlation across periods, which are critical to the comparative analysis conducted in this study.
@@
Figure 3: Spectral correlations: (a) target correlation structure, (b) empirical correlations from recorded motions, and (c) empirical correlations from synthetic motions. Contours represent correlation coefficients between spectral accelerations at periods $T_{1}$ and $T_{2}$. Panels (b) and (c) are computed using 2,000 ground motions in each dataset.
In addition, Figure 4 summarizes key ground motion intensity measures across the recorded and synthetic datasets, including peak ground acceleration (PGA), Arias intensity ($I_{a}$), and significant duration ($D_{5–95}$, defined as the time interval between 5% and 95% of Arias intensity accumulation). The mean values of each intensity measure are also reported. Notably, PGA and $I_{a}$ exhibit comparable distributions across the datasets, while $D_{5–95}$ is shorter for the synthetic motions, with smaller variability. It is also noted that within the recorded dataset, 418 motions are identified as near-fault events (defined by rupture distances less than 15km), and 118 motions exhibit characteristics commonly classified as pulse-like, according to the PEER database. The complete datasets are available through an open repository, as described in the Data Availability section.
Figure 4: Comparison of ground motion intensity measures between datasets: (a) PGA, (b) Arias intensity, and (c) significant duration. Histograms are based on 2,000 ground motions for each dataset.
2.4 Structural models
Two structural models are employed to investigate the influence of ground motion characteristics on seismic responses across distinct dynamic regimes: a 12-story medium-period building and a 110-story long-period tower. These models are selected to represent typical mid-rise and high-rise multi-degree-of-freedom (MDOF) structures, enabling a comparative evaluation of global and extreme response characteristics under synthetic and recorded ground motions.
@@ 2.4 Structural models
    Medium-period building: A 12-story shear-type MDOF system with a fundamental period of 1.09 seconds. The seismic response of this archetype is predominantly governed by low-mode dynamics and reflects typical behavior of mid-rise structural systems.
@@
    Long-period tower: A limiting case of 110-story shear-type MDOF system with a fundamental period of 9.69 seconds. Its seismic responses are significantly influenced by higher modes and low-frequency ground motion components***Although this case is presented as a limiting example of a very long-period structure, its structural system does not correspond to that of a realistic structure of this height..
[…]
The seismic responses of interest include peak interstory drift ratios (IDRs) and peak floor accelerations (PFAs), which are widely used in seismic performance evaluation. Nonlinear time-history analyses are conducted using OpenSees mckenna2011opensees .
Uncertainties in structural parameters $\boldsymbol{X}_{\mathbf{S}}$, including damping ratio, floor weight, yield force, story stiffness, and strength hardening ratio, are modeled using standard distribution models. Table 2 summarizes these distributions across four cases: Case 1 assumes uniform distributions, while Cases 2–4 employ lognormal and truncated Gaussian distributions with decreasing coefficients of variation (c.o.v). These uncertainties are consistently applied to both structural models in the comparative analysis. Figure 5 illustrates the structural models and the probability density functions (PDFs) of the parameters. Note that four test cases with different distribution models are considered to ensure the robustness of the comparative study conclusions.
Figure 5: Structural models: (a) medium-period building, (b) long-period tower, and (c) PDFs of uncertain structural parameters across four cases.
Table 2: Probabilistic distributions of uncertain structural parameters for the structural models.
@@ 3 Comparative analysis of seismic responses: synthetic versus recorded motions
This section presents a detailed comparative evaluation of seismic response metrics for synthetic and recorded ground motions. Five key metrics are analyzed: distributions, statistical moments, correlations, tail indices, and sensitivity indices. The first four metrics focus exclusively on aleatory uncertainties by considering ground motion variability, while structural parameters are fixed at their mean values, as specified in Case 1 of Table 2, ensuring that the analysis isolates the effects of ground motion variability on seismic responses. The fifth metric incorporates both ground motion variability and structural parameter uncertainty to evaluate their combined effect and relative contributions to seismic response variability. The analysis is conducted for both structural models, examining trends in global and extreme behaviors.